(x-6)/(x+5) = (x-4)/(3x+2)
Answer by edjones(8007)

(x-6)/(x+5) = (x-4)/(3x+2)
(x-6)(3x+2)=(x-4)(x+5)
3x^2-16x-12=x^2+x-20
2x^2-17x+8=0
8*2=16, two factors of 16 whose sum is -17 are -16 and -1
2x^2-x-16x+8=0
x(2x-1)-8(2x-1)=0 factoring by grouping.
(x-8)(2x-1)=0
x=8, x=1/2
.
Ed